# MetaCluster: Enabling Deep Compression of Kolmogorov-Arnold Network

### v1

@@ 5.2 HyperNetworks
Hypernetworks reduce trainable parameter counts by replacing task or instance-specific weights with a shared generator that predicts them on demand. The idea originates from early meta‐learning works such as Ba et al. (2016), where a small network generates classifier weights from context, and Bertinetto et al. (2016), which learns to emit a one‐shot tracker’s parameters conditioned on a single exemplar. Ha et al. (2016) then formalized HyperNetworks for CNNs and RNNs, showing that a compact hypernetwork can match the accuracy of a standard model while drastically cutting the number of directly optimized parameters. Later extensions use task embeddings to condition a single hypernetwork for multi‐task or multi‐objective weight generation (Savarese & Maire, 2019), (Navon et al., 2020) .
MetaKANs (Zhao et al., 2025) bring this paradigm to Kolmogorov–Arnold Networks by observing that the dominant cost in a KAN is storing the coefficients of every univariate activation. Instead of optimizing all coefficients directly, a small meta-learner maps per-activation embeddings to basis coefficients, capturing a shared rule for weight generation across activations (Zhao et al., 2025). In contrast, our method uses the meta‐learner only during training to impose a clusterable geometry on per‐edge coefficient embeddings. Then, at inference, we dispense with any hypernetwork, incurring zero runtime overhead while achieving even greater parameter efficiency.
@@ 5.3 Weight Sharing
Weight sharing compresses networks by restricting each layer’s parameters to a small set of shared centroids and storing a codebook, along with per-weight indices. Han et al. (2015) popularized the practice with their practical pipeline consisting of magnitude pruning and K-means weight sharing. Subsequent theory linked quantization error to loss curvature and leveraged a Hessian‑weighted K-means (Choi et al., 2016). Another line integrates clustering into training. For instance, differentiable K-means (DKM) optimizes centroids and assignments jointly with task loss, producing strong models in both vision and NLP (Cho et al., 2021). Building on prior work, our MetaCluster framework is the first to apply weight sharing to KANs successfully.
It is worth noting that, while weight-sharing is one avenue for model compression, bit-width quantization is an alternative avenue, which is largely orthogonal and complementary to the proposed approach. Applying quantization on top of MetaCluster (e.g., quantizing the code-book of centroids) has the potential to achieve a further reduction in the memory footprint.
@@ 6 Conclusion
We introduced MetaCluster, a compression framework that makes Kolmogorov–Arnold Networks practical at scale by introducing the novel combination of meta‑learned manifold shaping with weight sharing. A lightweight meta-learner maps low-dimensional embeddings to per-edge basis coefficients, constraining KAN activations to lie on a compact manifold that is amenable to clustering. We then apply K-means in coefficient space, replace per‑edge parameters with codebook centroids and compact indices, discard the meta‑learner, and briefly fine‑tune centroids to recover any loss. This design directly targets the dimensionality driver of KAN memory, yielding a storage advantage that grows with the number of coefficients per edge.
Across standard KANs and ConvKANs on MNIST, CIFAR‑10, and CIFAR‑100, MetaCluster achieves up to $80\times$ reduction in parameter storage without degrading accuracy. Visualizations and ablations demonstrate that manifold shaping is crucial for high-quality clustering in high dimensions, and that KANs benefit particularly from weight sharing compared to MLPs, as each centroid amortizes many coefficients.
@@ A.1 Hyperparameters
We report the complete set of hyperparameters for our experiments in Tables 4, 5, 6, and 7. For our fully-connected architecture experiments Table 5 reports the hyperparameters for our KAN variants and Table 5 reports the hyperparameters for our MetaKAN variants. Then for our convolutional architecture experiments Table 6 reports the hyperparameters for our KAN variants and Table 7 reports the hyperparameters for our MetaKAN variants.
@@
Learning rate for Metalearner (lr_h) $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-4}$

### v2

@@ 5.2 HyperNetworks
Hypernetworks reduce trainable parameter counts by replacing task or instance-specific weights with a shared generator that predicts them on demand. The idea originates in early meta-learning work such as Ba et al. (2016), which uses a small network to generate classifier weights from context, and Bertinetto et al. (2016), which learns to emit a one-shot tracker’s parameters conditioned on a single exemplar. Ha et al. (2016) then formalized Hypernetworks for CNNs and RNNs, showing that a compact hypernetwork can match the accuracy of a standard model while drastically cutting the number of directly optimized parameters. Later extensions use task embeddings to condition a single hypernetwork for multi-task or multi-objective weight generation (Savarese and Maire, 2019; Navon et al., 2020) .
MetaKANs (Zhao et al., 2025) bring this paradigm to KANs by observing that the dominant cost in a KAN is storing the coefficients of every univariate activation. Instead of optimizing all coefficients directly, a small meta-learner maps per-activation embeddings to basis coefficients, capturing a shared rule for weight generation across activations (Zhao et al., 2025). In contrast, our method uses the meta-learner only during training to impose a clusterable geometry on per-edge coefficient embeddings. Then, at inference, we dispense with any hypernetwork, incurring zero runtime overhead while achieving even greater parameter efficiency.
@@ 5.3 Weight Sharing
Weight sharing compresses networks by restricting each layer’s parameters to a small set of shared centroids and storing a codebook, along with per-weight indices. Scalar quantization approaches, popularized by Han et al. (2015) introduced a pipeline combining of magnitude pruning and K-means weight sharing. Subsequent theory linked quantization error to loss curvature and leveraged a Hessian-weighted K-means to cluster (Choi et al., 2016). Differentiable K-means (DKM) extended this idea by jointly optimizing centroids and assignments with task loss (Cho et al., 2021).
Beyond scalar quantization, vector quantization offers a natural extension by clustering weight vectors rather than individual scalars, which early work applied to improve compression–accuracy trade-offs (Gong et al., 2014). More recently, the rise of large language models has driven specialized vector quantization approaches, which leverage Hessian information and the inherent sparsity of LLMs to optimize vector-level assignments (Liu et al., 2024a; Egiazarian et al., 2024; Van Baalen et al., 2024; Tseng et al., 2024). We demonstrate in Section 4.1.1 that such Hessian information is not effective for KANs, a result that is in line with Hessian-aware quantization techniques reported in Fuad and Chen (2025). Building on vector quantization research, our MetaCluster framework is the first to apply weight-sharing principles to the KAN successfully.
It is worth noting that, while weight-sharing is one avenue for model compression, bit-width quantization is an alternative avenue, which is largely orthogonal and complementary to the proposed approach. Applying quantization on top of MetaCluster (e.g., quantizing the codebook of centroids) can further reduce the memory footprint.
@@ 6 Conclusion
We introduced MetaCluster, a compression framework that makes KANs practical at scale by combining meta-learned manifold shaping with weight sharing. A lightweight meta-learner maps low-dimensional embeddings to per-edge basis coefficients, thereby constraining KAN activations to lie on a compact manifold amenable to clustering. We then apply K-means in coefficient space, replace per-edge parameters with codebook centroids and compact indices, discard the meta-learner, and briefly fine-tune centroids to recover any loss. This design directly targets the dimensionality driver of KAN memory, yielding a storage advantage that grows with the number of coefficients per edge.
Evaluation shows that MetaCluster achieves up to $80\times$ reduction (up to $124.1\times$ on equation modeling) in parameter storage without degrading accuracy or performance. Visualizations and ablations demonstrate that manifold shaping is crucial for high-quality clustering in high dimensions, and KANs benefit particularly from weight sharing compared to MLPs, as each centroid amortizes many coefficients.
Impact Statement
This work advances machine learning systems research by making Kolmogorov-Arnold Networks significantly more memory-efficient without loss of accuracy. By reducing storage requirements, it may enable the use of expressive models in resource-constrained and scientific settings, with no ethical risks beyond those common to general-purpose model compression methods.
@@ A.1 Hyperparameters
A.1.1 Image Classification Fully-Connected Setup
Our fully-connected architecture follows Zhao et al. (2025) in that we stack two fully-connected KAN layers, with a 32-dimensional hidden state between them. Each KAN variant uses a SiLU activation (Elfwing et al., 2018). The meta-learner variants contained a meta-learner with a hidden dimension of 32. Concretely:
  • KAN: degree–3 B-splines, grid range [1,1][-1,1], grid size 55 (Liu et al., 2024b).
  • FastKAN: 88 RBFs over [2,2][-2,2] (Li, 2024).
  • GramKAN: degree–3 polynomial (Drokin, 2024).
We train with AdamW (Loshchilov and Hutter, 2017) and for at most 50 epochs, with early stopping (patience = 10). After the final epoch, we cluster the learned weights into 16 groups, then fine-tune for an additional 5 epochs.For our fully-connected architecture experiments Table 7 reports the full set of hyperparameters for our KAN variants and Table 8 reports the full set of hyperparameters for our MetaKAN variants.
@@ A.1.1 Image Classification Fully-Connected Setup
A.1.2 Image Classification Convolutional Setup
  • KANConv: degree–3 B-splines, grid range [3,3][-3,3], grid size 55 (Liu et al., 2024b).
  • FastKANConv: 88 RBFs on [3,3][-3,3] (Li, 2024).
  • GramKANConv: degree–3 polynomial (Drokin, 2024).
Training again uses AdamW for up to 150 epochs with early stopping (patience = 10). Final weights are clustered into 16 groups and fine-tuned for 5 epochs. For our convolutional architecture experiments Table 9 reports the full set of hyperparameters for our KAN variants and Table 10 reports the full set of hyperparameters for our MetaKAN variants.
@@ A.1.2 Image Classification Convolutional Setup
Learning rate for meta-learner (lr_h) $\times 10^{-4} 1×1$$\times 10^{-4}$ $1\times 10^{-4}$
@@
A.1.3 Equation Modeling Setup
We report our complete set of hyperparameter for our equation modeling experiments in Table 11.
Table 11: Hyperparameters for fully-connected KAN equation modeling experiments.
Hyperparameter KAN
Hidden dimension 32
Activation SiLU
Degree 3 (Spline)
Grid range [1,1][-1,1]
Grid size 5
Optimizer AdamW
Learning rate $1\times 10^{-4}$
Learning rate for finetuning (lr_c) $1\times 10^{-4}$
Batch size 256
Epochs 100
Clustered training epochs 10
Number of clusters 4
Domain of datapoints [-1,1]
Train Set samples 50,000
Test Set samples 20,000
@@ A.2.2 Meta-learner Embedding Size Extended
A.3 Visualizing Clusters
We verify the importance of MetaCluster qualitatively by plotting edge functions and the respective centroids associated with the first layer of FastKAN and MetaFastKAN. in Figures 4 and 5, respectively. From observing the functions associated with FastKAN and MetaFastKAN it is visually clear the functions associated with MetaFastKAN are much closer to the respective centroid than in FastKAN. Such a result demonstrates the importance using meta-learners to find a lower-dimensional functional space before clustering.
(a) The set of clusters showcasing the edge functions for each cluster on the first layer.
(b) The set of clusters showcasing the edge functions for each cluster on the second layer.
Figure 4: The edge functions for each cluster (in blue) and the centroid functions (in red) for the (a) first and (b) second layers of the fully-connected FastKAN models on Cifar-10.
(a) The set of clusters showcasing the edge functions for each cluster on the first layer.
(b) The set of clusters showcasing the edge functions for each cluster on the second layer.
Figure 5: The edge functions for each cluster (blue) and the centroid functions (red) for the (a) first and (b) second layers of the fully-connected MetaFastKAN models on Cifar-10.